OpenStudy (hnamowicz):

Which equation is a point slope form equation for line AB? x + 2y = 4 2x−y=22x−y=2 2x + y = 2 x−2y=4

OpenStudy (thekingofcr):

fist you need to get each equation in point slope intercept form

OpenStudy (thekingofcr):

can u do that?

OpenStudy (thekingofcr):

slop is 2 and y intercept is -2 look at the equations and see which one matches up once put into correct form

OpenStudy (hnamowicz):

i think its D

OpenStudy (thekingofcr):

x-2y=4 -x -2y=x+4 /-2 y=-1/2x-2 so no try again

OpenStudy (thekingofcr):

try c 2x+y=2 so subtract -2x from both sides y=-2x+2 and there ya go , so its c

OpenStudy (thekingofcr):

got it?

OpenStudy (hnamowicz):

i guess thanks

OpenStudy (hnamowicz):

it was 2x−y=2